OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±8) and asymptotes at y = ±4/ 3x

OpenStudy (anonymous):

y2/9 - x2/16=1

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

i think i got it

OpenStudy (anonymous):

y2/64-x2/36=1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

can u help me with this one

OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = -4 + 5 cos θ

OpenStudy (anonymous):

how would you check if it's symmetrical over the x-axis?

OpenStudy (anonymous):

the -4

OpenStudy (anonymous):

i think its symmtrical over the y axis?

OpenStudy (anonymous):

it's not symmetrical over the y-axis, but it's symmetrical over the x-axis

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

if f(-θ) = f(θ), then it's symmetrical over the x-axis if f(pi-θ) = f(θ), then it's symmetrical over the y-axis if you replace r with -r and you get the same equation, it's symmetrical over the origin